How do I solve -3x+2y=1
-5x-9y=-23

Answer by mananth(16946)

-3 x + 2 y = 1 .............1

-5 x -9 y = -23 .............2
Eliminate y
multiply (1)by 9
Multiply (2) by 2
-27 x 18 y = 9
-10 x -18 y = -46
Add the two equations
-37 x = -37
/ -37
x = 1
plug value of x in (1)
-3 x + 2 y = 1
-3 + 2 y = 1
2 y = 1 + 3
2 y = 4
y = 2